Solve each of the following systems by substitution.
5x - 2y=-5
y - 5x = 3

Answer by funmath(2873)

Solve each of the following systems by substitution.
L1) 5x - 2y=-5
L2) y - 5x = 3
Solve L2 for y (because y doesn't have a coefficient to worry about dividing) and substitute that in for y in L1 and solve L1 for x:
L2) y-5x+5x=5x+3 ----->>> y=5x+3
L1) 5x-2(5x+3)=-5
5x-10x-6=-5
-5x-6=-5
-5x-6+6=-5+6
-5x=1
-5x/-5=1/-5
x=-1/5
Plug this value into L2 and solve for y:
L2) y-5(-1/5)=3
y+1=3
y+1-1=3-1
y=2
The solution is (x,y)=(-1/5,2)
Happy Calculating!!!